# Bifurcations and canards in two coupled FitzHugh–Nagumo equations July 25, 2025

### v1

@@ Abstract.
We describe the slow-fast dynamics of two FitzHugh–Nagumo equations coupled symmetrically through the slow equations. We find an open set of parameter values for which the two equations synchronise, and another set with antisynchrony — where the solution of one equation is minus the solution of the other. We also obtain bistability — where these two types of solution coexist as attractors.
Canards are shown to give rise to mixed-mode oscillations. They also initiate small amplitude transient oscillations before the onset of large amplitude relaxation oscillations. We also discuss briefly the effect of asymmetric coupling, with periodic forcing of one of the equations by the other.
Keywords: FitzHugh–Nagumo; Symmetric Coupling; Bifurcations; Canards; Mixed Mode Oscillations.
@@ 1. Introduction
Although the FitzHugh–Nagumo equations (FHN) were created as a simplified model for nerve impulse, they have also been intensively studied for purely mathematical reasons. This is because they provide a very simple example of equations that show complex and varied dynamics.
Several different formulations of the FitzHugh–Nagumo equations [8, 9, 21] occur in the literature both as reaction-diffusion equations and as ordinary differential equations. A review can be found in Cebrián-Lacasa et al. [3]. Here we see it as the slow-fast system of ordinary differential equations
@@
Synchronisation of two coupled equations of Hodgkin-Huxley type (models for nerve impulse) has been studied by Labouriau & Rodrigues [20], but it does not cover FHN. Coupling two FHN has been addressed in many places in the literature. For instance, the effect of having the fast variable of a periodic solution of a FHN coupled to the fast equation of another FHN and thus forcing it has been explored numerically by several authors. It was described as experiments in an electrical circuit by Binczack et al. [1], through self-coupling by Desroches et al. [4] and also numerically by Hoff et al. [14]. Coupling FHN symmetrically through the fast equations has been also addressed both numerically and analytically in Pedersen et al. [22] and Kristiansen and Pedersen [16], numerically by Hoff et al. [14]. For both symmetric and asymmetric coupling through the fast equations, it is studied analytically by Campbell and Waite [2], numerically by Santana et al. [24] and with a delay in Saha and Freudel [23]. Two different coupling constants, one for the fast variables the other for the slow ones are explored by Kawato et al. [15] and by Krupa et al. in 2014 [17]. Other authors use the slow coordinate of a periodic solution of one FHN to force the fast coordinate of another, as in Doss-Bachelet et al. [5] and Krupa et al. in 2012 [18], the latter in a model with three time scales. Many of these authors explored the slow-fast structure of the equations focusing on several different aspects of the dynamics: synchrony, periodic and chaotic mixed-mode oscillations, canards, and bursts. We will discuss their findings in more detail in the final section of this article.
Here we couple two identical FHN symmetrically through the slow equations. To the best of our knowledge this has not been done before. The coupling may be understood as a diffusive term since it depends on the difference of the slow variables, that in the original formulation of FHN represents ion transport through a membrane. It allows us to compare the different types of coupling, as discussed in Section 6 below.
@@
After describing the model in Section 2 and its critical manifold in Section 3 we discuss conditions where the solution of the two equations synchronise in Section 4. Canards, giving rise to mixed-mode oscillations and small amplitude transients, are treated in Section 5. Finally our results are discussed in Section 6 comparing them to the findings by the above mentioned authors, with pointers to future work, followed by a brief excursion into the possibly chaotic mixed-mode oscillations that arise in one-directional coupling where one FHN is used for periodically forcing the other.
We have endeavoured to make a self contained exposition bringing together all related topics. We have drawn illustrative figures to make the paper easily readable. All figures in this article were created through numerical simulations conducted in Matlab, using integration functions such as ode15s or ode23s, except for a figure drawn by the authors.
@@ 2. The object of study
We study the dynamics of two FitzHugh–Nagumo equations coupled through the slow equations:
@@
This is a simplified model for the study of a network of two Hodgkin-Huxley neurones, as in two identical coupled nerve cells with dynamics given by $(x_{1},y_{1})$ and $(x_{2},y_{2})$. The parameter $k\neq 0$ is the coupling strength.
The parameters $b$ and $c$ modulate the dynamics of each cell as discussed in [10] and in [3]. We think of (2) as two coupled cells, each one with dynamics represented by $(x_{1},y_{1})$ and $(x_{2},y_{2})$. When $k=0$, the two cells are independent and the dynamics of (2) is characterised by the cartesian product of two FitzHugh–Nagumo equations. As described in [10], the dynamics of a single FitzHugh–Nagumo model is characterised by the existence of at least one and at most three equilibrium states, that may be either stable, unstable or saddles. For some parameter values there is also a periodic solution. This implies that for $k=0$ (no coupling) the dynamics of (2) is characterised by the existence of periodic orbits (cartesian product with one of the equilibria) and by a resonant hyperbolic torus foliated by periodic solutions of rotation number 1. Hopf bifurcations and canard explosions have been found in the particular cases $b=0$ (cf. [10, Example 6.1]) and $c=0$ (cf. [10, Example 6.2]). For $k\neq 0$ the dynamics of each oscillator interferes on the other. The normally hyperbolic torus (when it exists) persists for $k>0$ small.
System (2) is formulated as a slow-fast system with two fast and two slow equations that are, respectively, the equations for $\dot{x}_{i}$ and for $\dot{y}_{i}$, $i=1,2$. There is a huge literature on this type of system, we refer the reader to the book [19], whose treatment we follow.
Accordingly, we call $X=(x_{1},x_{2})$ the fast variables whose dynamics is governed by the fast equations $\varepsilon\dot{X}=F(X,Y)$ with $Y=(y_{1},y_{2})$ and
@@ 3. The critical manifold
An important concept in the dynamics of slow-fast systems is the critical manifold $C_{0}$ defined as the set of equilibria of the fast equations. In the singular case $\varepsilon=0$ this is the manifold where solutions lie. In the present case it is the surface
$C_{0}=\left\{(x_{1},x_{2},y_{1},y_{2})\in{\mathbf{R}}^{4}:\ y_{i}=\varphi(x_{i}),\ i=1,2\right\},$
that contains the fold lines where its projection into the slow variables $y_{1}$, $y_{2}$ is singular, i.e., where either $\varphi^{\prime}(x_{1})=0$ or $\varphi^{\prime}(x_{2})=0$. The set of fold lines is given by
@@
The degenerate points $(x_{1},x_{2},\varphi(x_{1}),\varphi(x_{2}))\in\Sigma$ with both $x_{1}=\pm x_{2}$ and $x_{1}=\pm 2/\sqrt{3}$ are double folds, that correspond to the transverse crossing of two fold lines in $C_{0}$.
@@
Let $v(\xi)$ be a smooth vector field in ${\mathbf{R}}^{n}$. A smooth manifold $M\subset{\mathbf{R}}^{n}$ invariant under the flow of $\dot{\xi}=v(\xi)$ is normally hyperbolic if there is a smooth splitting $\left.T{\mathbf{R}}^{n}\right|_{M}=TM\oplus N$ such that for all $\xi\in M$ the eigenvalues of $\left.Dv(\xi)\right|_{N}$ have non-zero real part.
@@
The sets $A$ and $S$ of the previous result may be written as the disjoint union described in the caption of Figure 1.
@@
The fast equations consist of a family of equations in the variables $x_{1}$ and $x_{2}$ parametrised by $Y=(y_{1},y_{2})$. In this context, for each $Y$ the manifold $C_{0}$ consists of isolated points, hence its tangent space is trivial and normal hyperbolicity reduces to the hyperbolicity of the equilibria of the fast equation. This is determined by the eigenvalues of the derivative of $F(X,Y)$ with respect to the fast variables $X$, i.e. by the matrix
$\dfrac{\partial F}{\partial X}{(x_{1},x_{2},y_{1},y_{2})}=\begin{pmatrix}\varphi^{\prime}(x_{1})&0\\ 0&\varphi^{\prime}(x_{2})\end{pmatrix}\qquad\mbox{where}\qquad\varphi^{\prime}(x)=4-3x^{2}\ .$
@@
A set $S\subset{\mathbf{R}}^{n}$ is locally flow-invariant under a vector field $W$ defined in ${\mathbf{R}}^{n}$ if for every point $x_{0}\in S$ there is a $t_{0}>0$ such that for every $t$ with $|t|<t_{0}$ the solution $x(t)$ of differential equation $\dot{x}=W(x)$ with $x(0)=x_{0}$ lies in $S$.
From Lemma 3.2 we obtain the following result:
Corollary 3.4.
For every $r\geq 2$ and for sufficiently small $\varepsilon>0$ there is a flow-invariant slow manifold $C_{\varepsilon}$ of class $C^{r}$ which is $\mathcal{O}(\varepsilon)$ close, in the Haussdorf topology, to the normally hyperbolic part of $C_{0}$ , where $\mathcal{O}$ stands for the usual Landau notation. Moreover, close to the set $A$ where $C_{0}$ attracts the fast flow, the manifold $C_{\varepsilon}$ is also locally attracting for the fast equations in (2) and there is a locally attracting and locally flow-invariant manifold $A_{\varepsilon}$.
@@
Since the critical manifold $C_{0}$ is normally hyperbolic everywhere except at the fold points in $\Sigma$, the result follows by Fenichel’s Theorem [7]. ∎
[…]
In special situations a trajectory may cross $\Sigma$ and remain in the unstable region of $C_{\varepsilon}$ for some time. Such a trajectory is called a canard and will be discussed in Section 5 below.
4. Synchrony
Equations (2) have the symmetry $\gamma(x_{1},x_{2},y_{1},y_{2})=(x_{2},x_{1},y_{2},y_{1})$, therefore the plane
$\left\{(x,x,y,y)\in{\mathbf{R}}^{4}\right\}=\operatorname{Fix}(\gamma)$
is flow-invariant. Solutions in this plane correspond to two synchronised cells: they behave like a single FitzHugh–Nagumo system. We refer to this plane as the synchrony plane. Equilibria in this plane satisfy both $y=\varphi(x)=4x-x^{3}$ and $x-b\varphi(x)-c=0$. Therefore in $\operatorname{Fix}(\gamma)$ there is always at least one equilibrium and there are at most three equilibria — details in Section 3 of [10]. By the Bézout Theorem there are at most 9 equilibria of (2) so there may be up to 3 symmetry related pairs of equilibria outside the synchrony plane.
The symmetry implies that the synchrony plane $\operatorname{Fix}(\gamma)$ is invariant under the flow of (2). The next result provides conditions for its intersection with the attracting part, $A$, of $C_{0}$ to be attracting.
Figure 2. Two singular solutions of (2) attracted to the synchrony plane shown in two different projections, parameters $b=0,c=-2$ and $k=1$. Conventions: trajectories with initial conditions at the brown and gray dots, slow part in brown/gray solid lines, fast part dotted, stable equilibrium (white dot), fold lines (red/black), double fold points (red dots), intersection with synchrony plane white line. (A) - Projection of $C_{0}$ into the subspace $\left(x_{1},x_{2},y_{1}\right)$ (purple). (B) - Same trajectories as in (A) projected into the plane $\left(x_{1},x_{2}\right)$.
Figure 3. Singular solution of (2) on the synchrony plane shown in two different projections, parameters $b=0=c$ and $k=1$. Conventions: trajectories with initial conditions at the brown dot, slow part in brown solid lines, fast part dotted, equilibria (white dots), fold lines (red/black) degenerate fold points (red dots), intersection with synchrony plane (white line). (A) - Projection $C_{0}$ into the subspace $\left(x_{1},x_{2},y_{1}\right)$ (purple). (B) - Same trajectories as in (A) projected into the plane $\left(x_{1},x_{2}\right)$.
Proposition 4.1.
The intersection $\operatorname{Fix}(\gamma)\cap A$ is normally hyperbolic if and only if $b+2k\geq 0$. Moreover, in this case it is locally attracting.
@@ 4. Synchrony
(5) $\left.D{\widehat{V}}(z_{1},z_{2},z_{3},z_{4})\right|_{(2x,2y,0,0)}=\begin{pmatrix}\varphi^{\prime}(x)&-1&0&0\\ 1&-b&0&0\\ 0&0&\varphi^{\prime}(x)&-1\\ 0&0&1&-(b+2k)\end{pmatrix}\qquad N_{\gamma}=\begin{pmatrix}\varphi^{\prime}(x)&-1\\ 1&-(b+2k)\end{pmatrix}$
where $N_{\gamma}$ is the component of $D\widehat{V}(Z)$ in the directions perpendicular to $\operatorname{Fix}(\gamma)$, evaluated at $Z=(2x,2y,0,0)\in\operatorname{Fix}(\gamma)$.
The synchrony plane is attracting if $\det(N_{\gamma})>0$ and $\operatorname{Tr}(N_{\gamma})<0$ where
$\det(N_{\gamma})=-(b+2k)\varphi^{\prime}(x)+1\quad\text{and}\quad\operatorname{Tr}(N_{\gamma})=\varphi^{\prime}(x)-(b+2k).$
As we saw in Section 3, points in the critical manifold attract the fast flow if $\varphi^{\prime}(x_{i})<0$ for $i=1,2$. So, if $b+2k\geq 0$ then $\operatorname{Fix}(\gamma)\cap A$ is normally hyperbolic and attracting, since $\det(N_{\gamma})>0$ and $\operatorname{Tr}(N_{\gamma})<0$.

### v3

# Synchrony and canards in two coupled FitzHugh–Nagumo equations July 25, 2025

@@ Abstract.
We describe the fast-slow dynamics of two FitzHugh–Nagumo equations coupled symmetrically through the slow equations. We use symmetry arguments to find a non-empty open set of parameter values for which the two equations synchronise, and another set with antisynchrony — where the solution of one equation is minus the solution of the other. By combining the dynamics within the synchrony and antisynchrony subspaces, we also obtain bistability — where these two types of solution coexist as hyperbolic attractors. They persist under small perturbation of the parameters. Canards are shown to give rise to mixed-mode oscillations. They also initiate small amplitude transient oscillations before the onset of large amplitude relaxation oscillations. We also discuss briefly the effect of asymmetric coupling, with periodic forcing of one of the equations by the other. We illustrate our results with numerical simulations.
Bruno F. F. Gonçalves
corresponding author
Centro de Matemática
Universidade do Porto
Rua do Campo Alegre, 687
4169-007 Porto
Portugal
Isabel S. Labouriau
Centro de Matemática
Universidade do Porto
Rua do Campo Alegre, 687
4169-007 Porto
Portugal
Alexandre A. P. Rodrigues
Lisbon School of Economics & Management
Centro de Matemática Aplicada
à Previsão e Decisão Económica
Rua do Quelhas 6
1200-781, Lisboa
Portugal
Keywords: FitzHugh–Nagumo; Symmetric Coupling; Synchrony; Canards; Mixed-mode Oscillations.
@@ 1. Introduction
The FitzHugh–Nagumo equations (FHN) have been intensively studied in their own right for purely mathematical reasons, as they provide a very simple example of equations that show complex and varied dynamics. This is the point of view of the present article. The equations have been introduced by FitzHugh [14, 15] as a simplified model for nerve impulse, but we do not address this aspect here. They have also been presented by Nagumo et al. [28] as a model for experiments in an electrical circuit.
Several different formulations of the FitzHugh–Nagumo equations [^1] Nagumo et al. [28] also quote it as Bonhoeffer–van der Pol equations. [14, 15, 28] occur in the literature both as reaction-diffusion equations and as ordinary differential equations. Reviews of the behaviour of a single FHN can be found for instance in Cebrián-Lacasa et al. [4] focusing on mathematical aspects of the dynamics and, with emphasis on neurobiology, in Dmitrichev et al. [9]. Here we see it as the following fast-slow system of ordinary differential equations
@@
Coupling two FHN has been addressed in many places in the literature. For instance, the effect of having the fast variable of a periodic solution of a FHN coupled to the fast equation of another FHN and thus forcing it has been explored numerically by several authors. It was described as experiments in an electrical circuit by Binczack et al. [2], through self-coupling by Desroches et al. [6] and also numerically by Hoff et al. [20]. Other authors have used the slow coordinate of a periodic solution of one FHN to force the fast coordinate of another, as in Doss-Bachelet et al. [10] and Krupa et al. [24], the latter in a model with three time scales. Dmitrichev et al. [9] have used the fast coordinate of one FHN to force the slow coordinate of another FHN.
Coupling FHN symmetrically through the fast equations has been addressed both numerically and analytically in Pedersen et al. [29] and Kristiansen and Pedersen [22]. The synchronisation of FHN with different parameter values coupled in this way was also treated analytically in Plotnikov and Fradkov [30]. Their dynamics was described numerically by Hoff et al. [20]. Both symmetric and asymmetric linear coupling of two FHN through the fast equations was studied analytically by Campbell and Waite [3] and numerically by Santana et al. [33]. This type of coupling, but with a delay was studied numerically in Saha and Freudel [32] and both numerically and experimentally in Ponomarenko et al. [31] and in Kulminskiy et al. [26] for synchronisation. Experimental and numerical results on both symmetric and asymmetric sigmoid coupling were described in Egorov et al. [11]. Two different coupling constants, one for the fast variables the other for the slow ones are explored by Kawato et al. [21] and by Krupa et al. in 2014 [23]. Synchronisation of two coupled equations of Hodgkin-Huxley type (models for nerve impulse) has been studied by Labouriau and Rodrigues [27], but it does not cover FHN.
Many of these authors explored the fast-slow structure of the equations focusing on several different aspects of the dynamics: synchrony, periodic and chaotic mixed-mode oscillations, canards, and bursts. We will discuss their findings in more detail in the final section of this article.
Here we couple two identical FHN symmetrically and linearly through the slow equations. To the best of our knowledge this has not been done before, probably because this type of coupling does not have a natural interpretation for nerve cells. However, this coupling could be realised, for instance, between two Nagumo’s circuits, as is suggested in Nagumo et al. [28, Figures 5 and 6], though the authors explore the connection in a different way. It allows us to compare the different types of coupling, as discussed at the end of this article.
@@
We explore analytically the two different time scales that arise in (1) for small values of $\varepsilon$ in order to obtain dynamical information on the coupled system. The fast-slow structure allows us to locate a region in phase space where all the recurring dynamics takes place. After describing the coupled equations in Section 2 we introduce some of the techniques of fast-slow systems in Section 3, in particular describing an important object in this type of study, the critical manifold. Conditions under which the solutions of the two equations synchronise are obtained in Section 4. A type of global dynamical behaviour called canards that gives rise to mixed-mode oscillations and small amplitude transients, is treated in Section 5. Finally our results are discussed in Section 6 comparing them to the findings by the above mentioned authors, with pointers to future work, followed by a brief excursion into the possibly chaotic mixed-mode oscillations that arise in one-directional coupling where one FHN is used for periodically forcing the other.
We have endeavoured to make a self contained exposition bringing together all related topics. We have provided illustrative figures and intuitive descriptions of the mathematical concepts and techniques to make the paper easily readable. For the reader not familiar with fast-slow systems we suggest reading our paper [16] where the method is described with the equations (1) as an example. All figures in this article were created through numerical simulations conducted in Matlab, using integration functions such as ode15s or ode23s, except for non-numerical figures drawn by the authors.
@@ 2. The object of study
In this section, we introduce the coupled system under analysis, as well as some terminology and notation that will be used throughout the article. More precisely, we study the dynamics of two FitzHugh–Nagumo equations coupled linearly through the slow equations:
@@
For $i\in\{1,2\}$, $\dot{x_{i}},\dot{y_{i}}$ represent the usual derivative with respect to $t$. We think of (2) as two coupled FHN systems, that we will abuse language calling cells, each one with dynamics represented by $(x_{1},y_{1})$ and $(x_{2},y_{2})$. The parameters $b$ and $c$ modulate the dynamics of each cell as discussed in [16] and in [4]. The parameter $k\neq 0$ is the coupling strength.
When $k=0$, the two cells are independent and the dynamics of (2) is characterised by the cartesian product of two FitzHugh–Nagumo equations. As described in [16], the dynamics of a single FitzHugh–Nagumo model is characterised by the existence of at least one and at most three equilibrium states, that may be either stable, unstable or saddles. For some parameter values there is also a type of periodic solution called relaxation oscillation whose speed has two times scales. This implies that for $k=0$ (no coupling) the dynamics of (2) is characterised by the existence of periodic orbits (cartesian product with one of the equilibria) and by a resonant normally hyperbolic [^2] See discussion of this concept in Section 3 below. torus foliated by periodic solutions of rotation number 1. Hopf bifurcations and canard explosions have been found in the particular cases $b=0$ (cf. [16, Example 6.1]) and $c=0$ (cf. [16, Example 6.2]). For $k\neq 0$ the dynamics of each oscillator interferes on the other. The normally hyperbolic torus (when it exists) persists for $k\neq 0$ small, as will be discussed in Section 3 below, where the concept of normal hyperbolicity is addressed.
System (2) is formulated as a fast-slow system with two fast and two slow equations that are, respectively, the equations for $\dot{x}_{i}$ and for $\dot{y}_{i}$, $i=1,2$. There is a huge literature on this type of system, we refer the reader to the book [25], whose treatment we follow.
A fast-slow system of differential equations is one in which some variables have their derivatives with larger magnitude than others. This leads to a system with two time scales. The general approach to this type of systems starts by grouping the variables in two disjoint sets: fast variables and slow variables. This separation is introduced in system (2) by the parameter $\varepsilon$. Accordingly, we call $X=(x_{1},x_{2})$ the fast variables whose dynamics is governed by the fast equations $\varepsilon\dot{X}=F(X,Y)$ with $Y=(y_{1},y_{2})$ and
@@ 3. The critical manifold
An important concept in the dynamics of fast-slow systems is the critical manifold $C_{0}$, defined as the set of equilibria of the fast equations (that in general are not equilibria of the full system). In the present case it is the surface
$C_{0}=\left\{(x_{1},x_{2},y_{1},y_{2})\in{\mathbf{R}}^{4}:\ y_{i}=\varphi(x_{i}),\ i=1,2\right\}.$
In the singular case $\varepsilon=0$ the equations (2) become a system of two differential equations in the variables $y_{i}$ subject to the constraint $(x_{1},x_{2},y_{1},y_{2})\in C_{0}$. The role of $C_{0}$ in the non singular case $\varepsilon>0$ will be discussed below, but before this we need to introduce another concept.
The critical manifold $C_{0}$ contains the fold lines (see Figure 3) where its projection into the slow variables $y_{1}$, $y_{2}$ is singular. In the present case this happens when either $\varphi^{\prime}(x_{1})=0$ or $\varphi^{\prime}(x_{2})=0$, hence the set of fold lines is given by
@@
The degenerate points $(x_{1},x_{2},\varphi(x_{1}),\varphi(x_{2}))\in\Sigma$ with both $x_{1}=\pm x_{2}$ and $x_{1}=\pm 2/\sqrt{3}$ are double folds, that correspond to the transverse crossing of two fold lines in $C_{0}$. They are illustrated as red points in Figures 2 and 3. The sets $C_{0}$ and $\Sigma$ have the structure of smooth manifolds.
Figure 1. Qualitative behaviour of a fast-slow system illustrated for the FitzHugh–Nagumo equations (1) with $b=c=0$ (the van der Pol equations). (A) - Solutions move rapidly to the critical manifold $C_{0}$ (thick red curve), then – as shown in (B) – they move slowly along the attracting part of $C_{0}$ (the solid thick red curve) under the slow equation until they reach a fold point, where they jump out of $C_{0}$ and move quickly to another attracting part of the critical manifold. The dashed curve in (B) is the repelling part of $C_{0}$ and in a two-dimensional system the folds are isolated points, in this case they are the local extremes of $y=\varphi(x)$.
An intuitive description of the dynamics in the non singular case $0<\varepsilon<\!\!<1$ is as follows. The coordinates $x_{1}(t)$ and $x_{2}(t)$ of a solution move a lot faster than the slow coordinates $y_{1}(t)$ and $y_{2}(t)$, so we may imagine that the $y_{i}(t)$ remain constant, while $(x_{1}(t),x_{2}(t))$ move fast to an attracting equilibrium of the fast equation, i.e., a point in $C_{0}$ where $\varphi^{\prime}(x_{i})<0$, as shown in Figure 1 (A) for a single FHN (1). Once in the critical manifold, the $x_{i}$ remain constant and the solutions evolve slowly in $C_{0}$, governed by the slow equations, and remain in $C_{0}$ until they reach a point in the fold lines. At these points one of the $\varphi^{\prime}(x_{i})$ is zero, so the equilibrium of the fast equation stops being attracting and the solution jumps out of $C_{0}$ as in Figure 1 (B). Away from $C_{0}$ the fast equation takes over and the solution moves quickly to another attracting equilibrium of the fast equation.
A rigorous version of the description above was initially obtained in 1952 by Tikhonov [35] (quoted, for instance, in [36]) and later extended by Fenichel [13]. Further extensions are described in [25]. In order to apply these results we need to define some concepts, valid for any vector field.
@@
Let $v(\xi)$ be a smooth vector field in ${\mathbf{R}}^{n}$, $S\subset{\mathbf{R}}^{n}$ and $M$ a smooth submanifold of ${\mathbf{R}}^{n}$.
  1. (1)
    The set $S$ is locally flow-invariant under $v$ if for every point $x_{0}\in S$ there is a $t_{0}>0$ such that for every $t$ with $|t|<t_{0}$ the solution $x(t)$ of $\dot{x}=v(x)$ with initial condition $x(0)=x_{0}$ lies in $S$.
  2. (2)
    The set $S$ is flow-invariant under $v$ if for every point $x_{0}\in S$ and every $t\in{\mathbf{R}}$, the solution $x(t)$ of $\dot{x}=v(x)$ with initial condition $x(0)=x_{0}$ lies in $S$.
  3. (3)
    The manifold $M\subset{\mathbf{R}}^{n}$ is normally hyperbolic if the component of the vector field $v$ transverse to $M$ is hyperbolic. More precisely, if there is a smooth splitting $\left.T{\mathbf{R}}^{n}\right|_{M}=TM\oplus N$ such that for all $\xi\in M$ the eigenvalues of $\left.Dv(\xi)\right|_{N}$ have non-zero real part where $T$ represents the tangent space.
Note that with these definitions a locally flow-invariant set may not be flow-invariant and a normally hyperbolic manifold is not hyperbolic. This may be confusing, but these are standard mathematical concepts, in line with the mathematicians’ practice of using adverbs. A result by Fenichel [13] shows that a manifold that is normally hyperbolic persists under small perturbations of the differential equations. We will be using this result throughout the article. In order to explore this feature we establish in the next result that the critical manifold has this property and classify the Lyapunov stability of the equilibria of the fast equations. Normal hyperbolicity will also be used in the study of synchrony in Section 4 below.
@@
The sets $A$ and $S$ of the previous result may be written as the disjoint union described in the caption of Figure 2. The letters $A$, $S$ and $R$ are suggestive of the stability character of the regions, $A$ for attracting, $S$ for saddle and $R$ for repelling.
@@
[…]
$\left.Dv(\xi)\right|_{N}=\dfrac{\partial F}{\partial X}{(x_{1},x_{2},y_{1},y_{2})}=\begin{pmatrix}\varphi^{\prime}(x_{1})&0\\ 0&\varphi^{\prime}(x_{2})\end{pmatrix}\qquad\mbox{where}\qquad\varphi^{\prime}(x)=4-3x^{2}\ .$
@@
The independence of the time scales when $\varepsilon=0$ makes the dynamics of (2) simpler to understand. The next result follows directly from Lemma 3.2 and says that the dynamics of the perturbed system (2) is similar to that of the singular system, with a deviation of $\mathcal{O}(\varepsilon)$, where $\mathcal{O}$ stands for the usual Landau notation. In that result the expression “close” refers to the Haussdorf topology.
Corollary 3.3.
With respect to (2), for every $r\geq 2$ and for sufficiently small $\varepsilon>0$, there is a locally flow-invariant manifold $C_{\varepsilon}$ of class $C^{r}$ which is $\mathcal{O}(\varepsilon)$ close to the normally hyperbolic part of $C_{0}$. Moreover, close to the set $A$ where $C_{0}$ attracts the fast flow, the manifold $C_{\varepsilon}$ is also locally attracting for the fast equations in (2) and there is a locally attracting and locally flow-invariant manifold $A_{\varepsilon}$.
Proof.
Since the critical manifold $C_{0}$ is normally hyperbolic everywhere except at the fold points in $\Sigma$, the result follows by Fenichel’s Theorem [13] that guarantees that for $0<\varepsilon\ll 1$, there exists a set $C_{\varepsilon}$ very close to $C_{0}$ exhibiting the same dynamical behaviour as $C_{0}$. ∎
The manifold $C_{\varepsilon}$ is called the slow manifold associated to (2).
[…]
In special situations a trajectory may cross $\Sigma$ and remain in the unstable region of $C_{\varepsilon}$ for some time. Such a trajectory is called a canard and will be discussed in Definition 5.1 below.
4. Synchrony and bistability
In this section we use symmetries to obtain conditions on the parameters of (2) under which its solutions are in synchrony: they behave like a single FitzHugh–Nagumo system. We show that the set of such synchronous solutions contains a subset that attracts all solutions that start near it. We also obtain conditions for antisynchrony where one of the “cells” in (2) is exactly minus the other, that also contains an attracting subset. We show that the two situations are compatible, giving rise to bistability in the equations, and we obtain conditions where these properties hold in an approximate sense when the two coupled equations are not identical. The proofs will involve normal hyperbolicity and rely on Fenichel’s Theorem [13]. We start by a rigorous description of the concepts.
@@ 4. Synchrony and bistability
Let $\xi(t)=\left(x_{1}(t),x_{2}(t),y_{1}(t),y_{2}(t)\right)$, $t\in{\mathbf{R}}$, be a solution of (2). Then we define:
  • $\xi(t)$ is in synchrony if we have $x_{1}(t)=x_{2}(t)$ and $y_{1}(t)=y_{2}(t)$ for all $t\in{\mathbf{R}}$; accordingly the plane ${\mathcal{S}}=\left\{(x,x,y,y):\quad x,y\in{\mathbf{R}}\right\}\subset{\mathbf{R}}^{4}$ is called the synchrony plane;
  • $\xi(t)$ is in antisynchrony if we have $x_{1}(t)=-x_{2}(t)$ and $y_{1}(t)=-y_{2}(t)$ for all $t\in{\mathbf{R}}$; accordingly the plane ${\mathcal{A}S}=\left\{(x,-x,y,-y):\quad x,y\in{\mathbf{R}}\right\}\subset{\mathbf{R}}^{4}$ is called the antisynchrony plane.
We start by addressing synchrony.
Theorem 4.2.
The synchrony plane ${\mathcal{S}}$ is flow-invariant under (2).
For small $\varepsilon>0$ the intersection ${\mathcal{S}}\cap A_{\varepsilon}$ of the synchrony plane with the attracting part, $A_{\varepsilon}$, of the slow manifold $C_{\varepsilon}$ is locally flow-invariant under (2). It is normally hyperbolic and locally attracting if $k>-b/2$.
@@
Equations (2) have the symmetry $\gamma(x_{1},x_{2},y_{1},y_{2})=(x_{2},x_{1},y_{2},y_{1})$, therefore the synchrony plane ${\mathcal{S}}$ is the set of fixed points of this symmetry ${\mathcal{S}}=\operatorname{Fix}(\gamma)$ and hence it is flow-invariant. The rest of the proof is divided in two lemmas stated and proved below.
In Lemma 4.3 we obtain conditions for the intersection of ${\mathcal{S}}=\operatorname{Fix}(\gamma)$ with the attracting part, $A$, of $C_{0}$ to be attracting. Then, in Lemma 4.4 we deal with the intersection with $A_{\varepsilon}$. ∎
Figure 3. Illustration of Lemma 4.3. Two singular solutions of (2) attracted to the synchrony plane shown in two different projections, parameters $b=0$, $c=-2$ and $k=1$. Conventions: trajectories with initial conditions at the brown and gray dots, slow part in brown/gray solid lines, fast part dotted, stable equilibrium (white dot), fold lines (red/black), double fold points (red dots), intersection with synchrony plane white line. (A) - Projection of $C_{0}$ into the subspace $\left(x_{1},x_{2},y_{1}\right)$ (purple) and trajectories. (B) - Same trajectories as in (A) projected into the plane $\left(x_{1},x_{2}\right)$.
Figure 4. Illustration of Theorem 4.2. Singular solution of (2) on the synchrony plane shown in two different projections, parameters $b=0=c$ and $k=1$. Conventions: trajectories with initial conditions at the brown dot, slow part in brown solid lines, fast part dotted, equilibria (white dots), fold lines (red/black), degenerate fold points (red dots), intersection with synchrony plane (white line). (A) - Projection $C_{0}$ into the subspace $\left(x_{1},x_{2},y_{1}\right)$ (purple). (B) - Same trajectories as in (A) projected into the plane $\left(x_{1},x_{2}\right)$.
In the next result we use properties of linear maps on ${\mathbf{R}}^{4}$ to analyse the stability. The expressions $\operatorname{tr}$ and $\det$ denote the usual trace and determinant operators.
Lemma 4.3.
The intersection $\operatorname{Fix}(\gamma)\cap A$ is normally hyperbolic if and only if $k\geq-b/2$. Moreover, in this case it is locally attracting.
@@
(5) $\left.D{\widehat{V}}(z_{1},z_{2},z_{3},z_{4})\right|_{(2x,2y,0,0)}=\begin{pmatrix}\varphi^{\prime}(x)/\varepsilon&-1/\varepsilon&0&0\\ 1&-b&0&0\\ 0&0&\varphi^{\prime}(x)/\varepsilon&-1/\varepsilon\\ 0&0&1&-(b+2k)\end{pmatrix}\qquad N_{\gamma}(x)=\begin{pmatrix}\varphi^{\prime}(x)/\varepsilon&-1/\varepsilon\\ 1&-(b+2k)\end{pmatrix}$
where $N_{\gamma}(x)$ is the component of $D\widehat{V}(Z)$ in the directions perpendicular to $\operatorname{Fix}(\gamma)$, evaluated at $Z=(2x,2y,0,0)\in\operatorname{Fix}(\gamma)$, with
$\det(N_{\gamma}(x))=-\dfrac{1}{\varepsilon}\left((b+2k)\varphi^{\prime}(x)-1\right)\quad\text{and}\quad\operatorname{tr}(N_{\gamma}(x))=\dfrac{\varphi^{\prime}(x)}{\varepsilon}-(b+2k).$
The synchrony plane is normally hyperbolic at $x\in\operatorname{Fix}(\gamma)$ if $\det(N_{\gamma}(x))\neq 0$ and $\operatorname{tr}(N_{\gamma}(x))\neq 0$. It is attracting at $x\in\operatorname{Fix}(\gamma)$ if $\det(N_{\gamma}(x))>0$ and $\operatorname{tr}(N_{\gamma}(x))<0$.
As we saw in Section 3, points in the critical manifold attract the fast flow if $\varphi^{\prime}(x_{i})<0$ for $i=1,2$. So, if $k\geq-b/2$ then $\operatorname{Fix}(\gamma)\cap A$ is normally hyperbolic and attracting, since $\det(N_{\gamma}(x))>0$ and $\operatorname{tr}(N_{\gamma}(x))<0$.